A sealed metal container contains a gas at37.0 ?Cand 1.00 atm.


To what temperature must the gas be heated for the pressure to double to 2.00 atm? (Ignore expansion of the container.)
Answer #1

pv = nrt

p/t = contant

1/ 310 = k

2/T = k

T = 2*310 = 620K

= 620-273 = 347 oC